Im finding X
Triangle FGH is an equilateral triangle with FG=X+5, GH=3X-9, andFH=2X-2 what is X equal

Answer by Simnepi(216)

Because it is an equilateral triangle all the sides are the same, that is,
FG = GH = FH
Or
x+5 = 3x-9 = 2x-2
Let's have a look at just one pair of these sides, FH and GH.
2x-2 = 3x-9
Looking at these we can see that GH has one more x than FH but is also 7 less in the number part of the expression.
Because these sides are equal, we have not increased or decreased the side, we can say that x must be the same as 7. Therefore
x = 7
To check:
Work each expression out and see if they are the same.
x+5 = 7+5 = 12
3x-9 = 21-9 = 12 and
2x-2 = 14-2 = 12.
Yay! Our answer must be correct! so x = 7